# Adaptive Super-Resolution Imaging Without Prior Knowledge Using a Programmable Spatial-Mode Sorter

### v1

@@
Abstract
We consider an imaging system tasked with estimating the angular distance between two incoherently-emitting sub-Rayleigh-separated point sources, without any prior knowledge of the centroid of the constellation and with a fixed collected-photon budget. It was shown theoretically that splitting the optical recording time into two stages—focal-plane direct imaging to obtain a pre-estimate of the centroid, and using that estimate to center a spatial-mode sorter followed by photon detection of the sorted modes—can achieve 10 to 100 times lower mean squared error in estimating the separation [1]. In this paper, we demonstrate this in proof-of-concept, using a programmable mode sorter we have built using multi-plane light conversion (MPLC) using a reflective spatial-light modulator (SLM) in an emulated experiment where we use a single coherent source to characterize the MPLC to electronically piece together the signature from two closely-separated quasi-monochromatic incoherent emitters.
journal: opticajournal
@@ 1 Introduction
For both adaptive and non-adaptive imaging, directly passing light from the scene through conventional optics like lenses and mirrors before reaching a focal-plane detector array constitutes the most widespread method. While effective for many purposes, this "direct imaging" approach encounters a resolution constraint often referred to as the Rayleigh limit ($\theta_{\rm{res}}$) [6], an angular distance scale beyond which the features of the scene become exceedingly hard to resolve. The limitation in resolution stems from diffraction effects arising due to the finite aperture of the imaging system, as illustrated in Fig.1. For a telescope-like system with a circular aperture, this limit is defined as follows:
@@
While sub-diffraction imaging can be done in the near field [7], with controllable fluorescent probes [8, 9, 10, 11], and via structrured optical excitation [12], fully passive optical super-resolution was made possible by a recent class of imaging receivers known as spatial mode demultiplexing (SPADE) [13]. SPADE uses physical linear optical processes such as multi-path interferometry [14] or holography [15] to couple specific orthogonal spatial modes of the incoming optical field to corresponding intensity detectors. References present various experimental demonstrations showcasing resolution enhancement below the Rayleigh limit in the transverse distance estimation problem through the utilization of spatial mode demultiplexing to classify Hermite-Gauss (HG) modes, prior to detection. Recently, multi-plane light conversion (MPLC), which uses a succession of phase masks to transform the optical wavefront [16], has emerged as the most promising SPADE implementation. Notably, a commercially available MPLC device that demultiplexes a fixed set of 2D spatial modes (Hermite Gauss modes) has spurred experimental progress for several recently demonstrated sub-diffraction imaging tasks [17, 18, 19, 20].
Neither SPADE nor MPLC are constrained to a particular set of spatial modes, and a device that can update operational spatial mode basis would be a perfect fit for adaptive super-resolution imaging. Several theoretical works have proposed using mid-acquisition adaptation to extend passive super-resolution techniques to more complex imaging tasks [1, 21, 22, 23]. Experimental progress in adaptive mode sorting is limited to date, with some systems demonstrating limited adaptability [24]. A fully adaptive MPLC implementation could perform highly complex imaging tasks in the presence of nuisance parameters.
Our experiment attempts to complete a more complicated task, demonstrating resolution enhancement below the Rayleigh limit in the transverse distance estimation problem without prior knowledge of the locations of the two incoherent sources or the centroid of the scene [1]. In addition, we employ a reconfigurable mode sorter, potentially applicable in more challenging localization tasks in subsequent iterations [23].
2 Main Results
@@
Our experiment primarily aims to estimate the separation ($\theta$) between two identical incoherent point sources with an unknown centroid, illustrated in Figure 1. To accomplish this, we adopted the strategy that was first hinted in [13] and fully outlined in [1], which suggests dividing photon resources between direct detection for centroid estimation and binary mode sorting for separation estimation. Direct detection has the advantage of shift invariance, allowing it to identify the location of the source pair over a wide field of view, and also is close to an optimal measurement of the object centroid [25]. Our second-stage SPADE measurement has been shown to achieve the resolution quantum limit for estimating small separations between point sources when the mode sorter is nearly perfectly aligned [13].
@@
The principal findings, expounded upon in the subsequent results section, reveal notable enhancements. Specifically, our hybrid approach demonstrates a reduction in variance of separation estimation for separations of up to 0.3 Rayleigh units, as well as a decrease in mean squared error of separation estimation for separations of up to 0.2 Rayleigh units when contrasted with direct imaging. The estimation of separations is executed using maximum likelihood estimators applied to both the direct imaging data and the sorted data.
The findings outlined in this paper serve as a proof of concept. It is important to note that the experimental results are partly emulated, given that we did not employ two distinct incoherent sources simultaneously. Instead, data acquisition was conducted sequentially from individual sources and subsequently merged computationally. Additionally, we took compensatory measures to account for losses attributable to our mode sorter, as detailed in the supplementary materials. This ensured parity in photon count between the two compared methods: direct imaging and the hybrid approach of half direct imaging and half mode sorting.
@@ 3 System Overview
Figure 2: (a) sample phase masks that would be displayed on the SLM, where SLM 1 acts as the object under test and the following three masks act as the MPLC masks. (b) Experimental setup.
@@ 3.1 System walk through
Our system consists of 4, main, subsystems:
@@
Illumination: Our 532 nm laser source (Edmund Optics 10 mW DPSS) first goes through a linear polarizer in order to align its polarization to the polarization required by the SLM. A spatial filter is then used to eliminate the higher spatial frequencies of the beam and produce a clean, uniform, Gaussian beam. The filtered beam is then collimated and directed to the SLM.
@@
SLM: In our system we use a liquid crystal based SLM (Holoeye PLUTO-2.1-VIS-016), which will cause aberrations in the system if not accounted for (See supplementary material). As can be seen in Fig.2(a), four square regions are displayed on the SLM. The first region is used to encode a phase-modulated object, and the other three are used as phase masks created by our algorithm, which was based on [16], forming the MPLC-based optical transformation that will ideally direct each orthogonal input mode to a respective, pre-set, location on the detector. On each square, we applied a blazed grating and used the first diffraction order rather than the zeroth order. The zeroth order is noisier due to the reflection of unmodulated light, while the first diffraction order is lossier. We use the latter as our focus is to minimize the crosstalk in mode sorting. Note that due to the SLM’s programmability and the fact that our MPLC- based code works for every orthogonal mode set, we achieve a reconfigurable mode sorter.
@@
4f system: As seen in Fig.2(b), the light goes through a 4f system right after the object is encoded at the first bounce from the SLM. The reason that we need the 4f system is to blur our phase-encoded object. By placing a pinhole (PH2 in Fig.2(b)) at the Fourier plane of the 4f system we effectively create a low pass filter, and by changing the size of PH2 (in our experiment we used a 200 microns pinhole, we can change the width of the PSF. This allows us to experiment with super-resolution imaging even though our object itself is not small enough. Due to the nature of the 4f system, the conjugate planes at the start and the end of the 4f system will be identical (if we take PH2 out) and therefore the phase-encoded object is going to be presented to the remaining SLM planes for processing.
@@
Detection: As shown in Fig.2(b), we use a beam splitter (BS) to be able to access both a direct image and the mode demultiplexed detection data. Note that only one of the outputs is going to reach the camera at a time and this will be decided by the flip mirror (M8 in Fig.2(b)). Each of these outputs is detected using the photon counting feature of an electron-multiplying charge-coupled device (EMCCD) camera (Andor iXon 897).
@@
A more detailed system walk-through can be found in [26].
@@ 3.2 Zernike Modes Sorting Characterization
The ideal modes to sort for our experiment’s purposes are the Zernike (or airy disc) mode family, since it is the PSF-adapted mode for hard circular aperture [27]. To ensure the feasibility of the experiment we tested how well our mode sorter sorts the Zernike modes of interest. To conduct that test we displayed a point source as the object (SLM 1 in Fig.2a) and swept it across the whole scene. As can be seen in Fig.3, the experimental data from the point source sweep closely matches the theoretical results, especially when the deviation off-axis is small. When the object is exactly on axis (where theoretically we expect all the photons to be in the Zernike00 mode) we saw only 0.89% cross-talk (definition of cross-talk further discussed in the supplementary material).
Note that while in [1] binary mode sorting is discussed, in Fig.3 we can see sorting of 3 modes. In our experiment, we still tried to imitate the binary mode sorting case. However, as the separation between the point sources increased we saw more contribution from higher order mode (as expected). To mitigate this contribution and ensure that it does not corrupt our data we decided to sort 3 modes (Zernike00, Zernike01, and Zernike02). The objective of sorting the third mode (Zernike02) was to ensure that the photons that live in the higher-order modes (especially when the separation increases) are directed there rather than to the other two spots which would corrupt the data. The other two spots (Zernike00 and Zernike01) were then used to evaluate the binary mode sorter’s performance in super-resolution imaging.
Figure 3: Calibration plot of theoretical (solid lines) and experimental (circles) intensities found in each mode with respect to the distance of the point source away from the known center. Amplitudes of the three orthogonal spatial modes that are sorted by the MPLC are shown on the left. Below the plot are the raw experimental images with the regions designated for each mode shown in circles, where each image corresponds to the indicated distance away from the system’s axis. Note: we report crosstalk value of 0.9831 between the fundamental mode and the higher order modes when the point source transition=0.
@@ 4.1 Experiment Overview
Our experiment is divided into 2 different stages, and its objective is to prove the capability of our mode sorting system to achieve super-resolution capabilities with relaxed prior information. More specifically, we wanted to exceed the Rayleigh limit [6] when estimating the separation between two point sources without knowing their centroid. The way in which we conducted our 2-stage experiment can be seen in [1]. To summarize, first, direct imaging data was acquired, from this data the centroid of the two point sources can be approximated (Section 4A).
Once the centroid is approximated, our MPLC-based algorithm creates the appropriate phase masks for the found centroid to perform the mode sorting. This generation of special phase masks essentially aligns our system to the scene and eliminates any need for physical manipulations to the system. Finally, using these specially created phase masks, mode sorting is performed. From this data (sorted data), we estimate the separation between the two point sources using maximum likelihood estimation as part of the data processing. In this experiment, we tested 12 different separations (from 0.1 Rayleigh units of separation to 1.2 Rayleigh units of separation in increments of 0.1 Rayleigh units) and had 200 data points for each separation. Each data point had a randomly selected centroid within the field of view of our system.
All the data in this experiment was taken using the photon counting mode in our EMCCD and the total number of photon arrivals at each data point (which is a combination of two images) was aimed to be 10,000. The laser power going into the system was calibrated accordingly while ensuring a low likelihood of multiple photons events. (See supplementary materials section D). Note: We conducted the experiment using a laser (coherent source), while for the experiment to correctly correlate the theory of super-resolution for two incoherent point sources, the light from the two point sources cannot interfere with each other. We therefore acquired data from only one point source at a time and then combined the two images computationally.

### v2

@@
Adaptive Super-Resolution Imaging Without Prior Knowledge Using a Programmable Spatial-Mode Sorter
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
We consider an imaging system tasked with estimating the angular distance between two incoherently-emitting, identically bright, sub-Rayleigh-separated point sources, without any prior knowledge of the centroid or the constellation and with a fixed collected-photon budget. It was shown theoretically that splitting the optical recording time into two stages—focal-plane direct imaging to obtain a pre-estimate of the centroid, and using that estimate to center a spatial-mode sorter followed by photon detection of the sorted modes—can achieve lower mean squared error in estimating the separation [1]. In this paper, we demonstrate this in a proof-of-concept, using a programmable mode sorter we have built using multi-plane light conversion (MPLC) using a reflective spatial-light modulator (SLM) in an emulated experiment where we use a single coherent source to characterize the MPLC to electronically piece together the signature from two closely-separated quasi-monochromatic incoherent emitters. We show an improvement in estimator variance when compared to direct imaging, in good agreement with simulations.
@@ 1 Introduction
For both adaptive and non-adaptive imaging, directly passing light from the scene through conventional optics like lenses and mirrors before reaching a focal-plane detector array constitutes the most widespread method. While effective for many purposes, this "direct imaging" approach encounters a resolution constraint often referred to as the Rayleigh limit [6], an angular distance scale beyond which the features of the scene become exceedingly hard to resolve. The limitation in resolution stems from diffraction effects arising due to the finite aperture of the imaging system, as illustrated in Fig.1. For a telescope-like system with a circular aperture, the Rayleigh limit sets a characteristic length scale
@@
While sub-diffraction imaging can be done in the near field [7], with controllable fluorescent probes [8, 9, 10, 11], and via structrured optical excitation [12], fully passive optical super-resolution was made possible by a recent class of imaging receivers known as spatial mode demultiplexing (SPADE) [13]. SPADE uses physical linear optical processes such as multi-path interferometry [14] or holography [15] to couple specific orthogonal spatial modes of the incoming optical field to corresponding intensity detectors. References present various experimental demonstrations showcasing resolution enhancement below the Rayleigh limit in the transverse distance estimation problem through the utilization of spatial mode demultiplexing to classify Hermite-Gauss (HG) modes prior to detection [16, 17, 18, 19]. Recently, multi-plane light conversion (MPLC) [20, 21, 22], which uses a succession of phase masks to transform the optical wavefront (Fig. 2(a)), has emerged as arguably the most promising SPADE implementation[16, 17, 18, 19]. Notably, a commercially available MPLC device that demultiplexes a fixed set of 2D spatial modes (Hermite Gauss modes) has spurred experimental progress for several recently demonstrated sub-diffraction imaging tasks [19, 18, 23, 17].
Neither SPADE nor MPLC are constrained to a particular set of spatial modes, and a device that can update operational spatial mode basis would be a perfect fit for adaptive super-resolution imaging. Several theoretical works have proposed using mid-acquisition adaptation to extend passive super-resolution techniques to more complex imaging tasks [1, 24, 25, 26]. Experimental progress in adaptive mode sorting is limited to date, with some systems demonstrating limited adaptability [27]. A fully adaptive MPLC implementation could perform highly complex imaging tasks in the presence of nuisance parameters[26, 28].
Our experiment attempts to complete a more complicated task, demonstrating resolution enhancement below the Rayleigh limit in the transverse distance estimation problem without prior knowledge of the locations of the two identically bright, incoherent sources or the centroid of the scene [1]. In addition, we employ a spatial light modulator (SLM) to implement a reconfigurable mode sorter, enabling an adaptive two-stage measurement procedure that uses the results of a first-stage direct imaging measurement to algorithmically align the second-stage SPADE device. Using an adaptive optical device (e.g., an SLM) to replace an optomechancial pointing apparatus has significant potential advantages for real-world imaging systems including size, weight and power (SWaP) considerations [29], alignment latency (60 Hz refresh rate for the SLM in our experiment), and pointing stability, especially considering the rapid development of compact, tunable metasurfaces for optical modulation [30].
2 Main Results
@@
Our experiment aims to estimate the separation ($\theta$) between two identical incoherent point sources with an unknown centroid ($\phi$), illustrated in Figure 1. To accomplish this, we adopted the strategy that was first hinted in [13] and fully outlined in [1], which suggests dividing photon resources between direct detection for centroid estimation and binary mode sorting for separation estimation. Direct detection has the advantage of shift invariance, allowing it to identify the location of the source pair over a wide field of view, and also is close to an optimal measurement of the object centroid [31]. Our second-stage SPADE measurement has been shown to achieve the resolution quantum limit for estimating small separations between point sources when the mode sorter is nearly perfectly aligned [13]. Theoretically, the optimal ratio for dividing the photon resources is separation ($\theta$) dependent and can be determined adaptively, which was demonstrated via simulations in [1]. In our experiment, we divided the photons equally between the two stages for simplicity, while adaptively utilizing the centroid estimate from first-stage direct imaging to align the second-stage mode sorter.
@@
The principal findings, expounded upon in the subsequent results section, reveal notable enhancements. Specifically, our hybrid approach demonstrates a reduction in variance of separation estimation for separations of up to a factor of 0.3 of the Rayleigh criterion, as well as a decrease in mean squared error of separation estimation for separations of up to textcolorreda factor of 0.2 of the Rayleigh criterion when contrasted with direct imaging. The estimation of separations is executed using maximum likelihood estimators applied to both the direct imaging data and the sorted data.
It is important to note that the experimental results are partly emulated, both in that we did not employ two distinct incoherent sources simultaneously but computationally combined data that were acquired sequentially from individual point sources, and also due to the fact that we did not implement real-time adaptation but performed the two stages of our measurement asynchronously. Nevertheless, the findings outlined in this paper serve as a proof of concept confirming the experimental viability of adaptive spatial mode sorting for sub-diffractive imaging.
@@ 3 System Overview
Figure 2: (a) Sample phase masks that would be displayed on the SLM, where SLM 1 acts as the object under test and the following three masks act as the MPLC masks. (b) Experimental setup. L532: 532 nm laser. HWP: half wave plate. PBS: polarized beam splitter. BB: beam block. M: mirror. SP-FILTER: spatial filter for single-mode beam processing. CL: cylindrical lens. SLM: spatial light modulator. DM: D-shaped pickoff mirror. 4fL: relay lens for 4f system. PH: pinhole. BS: non-polarzing beamsplitter. RP: right-angle prism. M8: flip mirror. EMCCD: electron-multiplying charge-coupled device camera.
@@ 3.1 System walkthrough
Our experimental setup consists of 4 main subsystems:
@@
Illumination: Our 532 nm laser source (Edmund Optics 10 mW DPSS) first goes through a linear polarizer in order to align its polarization to the polarization required by the SLM. A spatial filter is then used to eliminate the higher spatial frequencies of the beam and produce a clean, expanded, and uniform, Gaussian beam. The filtered and expanded beam is then collimated and directed to the SLM. Our goal is to use the expanded gaussian beam to uniformly illuminate the first SLM region (trying to get as close to a flat top). This will yield a scene with uniform intensity distribution.
@@
SLM: In our system we use a liquid crystal based SLM (Holoeye PLUTO-2.1-VIS-016, 8 $\mu$m pixel pitch), which applies a polarization-sensitive phase modulation to incident light at 532 nm in response to pixel-dependent, digitally driven voltage inputs. As can be seen in Fig.2(a), four square regions (300x300 pixels) are displayed on the SLM. The first region is used to encode an intensity-modulated object, and the other three are used as phase masks created by our algorithm, which was based on [20], forming the MPLC-based optical transformation that will ideally direct each orthogonal input mode to a respective, pre-set, location on the detector. On each square, we applied a blazed grating (period of 4.2 pixels on SLM region 1, period of 8 pixels on SLM regions 2-4) and coupled the first diffraction order of each blazed grating into the beampath rather than the zeroth order. The zeroth order is noisier due to the reflection of unmodulated light, while the first diffraction order is lossier. We use the latter as our focus is to minimize the crosstalk in mode sorting. The blazed grating on the first SLM region was used to generate binary-encoded intensity modulated objects by only coupling light into the continuing beampath from the portions of the SLM region containing the blazed grating. The blazed grating on the second, third and fourth SLM regions were statically displayed across the whole region over top of the MPLC phase masks. We additionally applied a correction phase mask over top of the entire SLM display to correct the systematic static wavefront aberration we characterized for our SLM device (see Supplementary Material Section 2). Note that due to the SLM’s programmability and the fact that our MPLC algorithm works for any orthogonal mode set, we achieve a reconfigurable mode sorter (see Supplementary Material Section 1 for sorting performance using different orthogonal mode sets).
@@
4f system: The light next goes through a 4f system consisting of two plano-convex lenses of focal length $f=200$ mm right after the object is encoded at the first bounce from the SLM. The purpose of the 4f system is to introduce diffractive blur on the intensity-encoded object by placing a pinhole (PH2 in Fig.2(b)) at the Fourier plane of the 4f system to effectively create a low pass spatial filter. By choosing the diameter of PH2 (in our experiment we used a pinhole with diameter $D=200\mu$m), we can change the width of the PSF. This allows us to use our tabletop optical setup to conduct a super-resolution imaging experiment with a characteristic PSF width of $\theta_{\rm res}=1.22*532\textrm{ nm}*200\textrm{ mm}/200\,\mu\textrm{m}=649\,\mu\textrm{m}$, equal to 81.125 SLM pixels or 40.56 camera pixels. The conjugate planes at the start and the end of the 4f system are coupled to the first and second SLM regions, respectively, allowing the intensity-encoded object to be presented to the remaining SLM planes for MPLC processing.
MPLC: Our 3-plane MPLC uses successive reflections off of the last three SLM regions. Two right angle prisms (RP) placed at a distance of 100 mm redirect the light back to the next SLM region. Prisms were used instead of a single mirror reflector in order to achieve an optical design that separately accesses SLM region 1 for object modulation. The MPLC algorithm from [20] been modified to account for this geometry.
@@
Detection: Finally, we use a beam splitter (BS) to be able to access both a direct image and the mode demultiplexed detection data. Note that only one of the outputs is going to reach the camera at a time and this will be decided by the flip mirror (M8 in Fig.2(b)). The camera is positioned on the table such that the detector plane lies at a conjugate image plane to the the first SLM region (with unit magnification) when M8 is flipped down and at the programmed output plane of the MPLC when M8 is flipped up. Each of these outputs is detected using the photon counting feature of an electron-multiplying charge-coupled device (EMCCD) camera (Andor iXon 897, 16 $\mu$m pixel pitch). See Section 4 of the Supplementary Material for details on photon counting characterization.
@@
A more detailed system walk-through of our experimental setup can be found in [32].
@@ 3.2 Zernike Modes Sorting Characterization
Since we used a circular pinhole as our aperture, our imaging system has a 2D Airy disk PSF. The associated coherent PSF has the functional form
$\psi(x,y)=\frac{J_{1}\left(1.22\pi r/\theta_{\rm res}\right)}{0.61\pi r/\theta_{\rm res}}$ (2)
where $J_{1}(\cdot)$ is the Bessel function of the first kind and where
$r=\sqrt{x^{2}+y^{2}}.$ (3)
The characteristic resolution scale $\theta_{\rm res}$ [Eq. (1)] is defined to match the radial distance to the first zero in the Airy ring diffraction pattern.
[…]
To ensure the feasibility of the experiment we tested how well our mode sorter sorts the PAD modes of interest. To conduct that test we displayed a blazed grating on only a 20x20 pixel square on the first SLM region (SLM 1 in Fig. 2a) to approximate a point source and swept it across the whole scene. We chose a 20x20 pixel size for our point source because it was empirically the smallest region we could designate for a blazed grating before a sharp dropoff in the reflectivity into the first blazed grating order. In Fig. 3 we plot the experimentally characterized mode intensity distribution as a function of point source spatial translation, where the term "Rayleigh units" means that the x-axis is scaled by the distance to the first zero in the Airy disk. The experimental data from the point source sweep closely matches the theoretical expectation, which was numerically calculated from the overlap of the shifted Airy disk with especially when the deviation off-axis is small. When the object is exactly on axis (where theoretically we expect all the photons to be in the PAD00 mode) we saw only 1.69% cross-talk (definition of cross-talk further discussed in the Supplementary Material Section 1).
Our experiment sought to imitate the binary mode sorting case discussed in [1]. However, as the separation between the point sources increased we saw more contribution from higher order modes (as expected) and any light from unsorted modes will randomly spread across the camera after passing through the MPLC, contributing noise to the measurement. To mitigate this contribution we decided to sort 3 modes (PAD00, PAD01, and PAD02, see left side of Fig. 3). The objective of sorting the third mode (PAD02) was to ensure that the photons that live in the higher-order modes (especially when the separation increases) are directed there rather than to the other two spots which would corrupt the data (see bottom of Fig. 3). The other two spots (PAD00 and PAD01) were then used to evaluate the binary mode sorter’s performance in super-resolution imaging. The decision to sort only 3 modes is based on the fact that sorting more modes hurts the performance of the mode sorter and produces more crosstalk (see Supplementary Material, Section 1).
Figure 3: Calibration plot of theoretical (solid lines) and experimental (circles) intensities found in each mode with respect to the distance of the point source away from the known center The step size was discretized by the SLM pixels (each pixel equate to $\approx$ 0.0125 Rayleigh units) . Amplitudes of the three orthogonal spatial modes that are sorted by the MPLC are shown on the left. Below the plot are the raw experimental images with the regions designated for each mode shown in circles, where each image corresponds to the indicated distance away from the system’s axis. Note: we report crosstalk value of 1.69% between the fundamental mode and the higher order modes when the point source translation=0.
@@ 4.1 Experiment Overview
The objective of our experiment is to prove the capability of our mode sorting system to achieve super-resolution parameter estimation with relaxed prior information. More specifically, we sought to demonstrate the theoretical conclusion of Ref. [1]: that adaptive spatial mode sorting can exceed the two-point-source separation estimation precision of direct imaging without knowing the object centroid a priori. We conducted our 2-stage experiment using the following procedure, of which we performed 200 independent trials for each of twelve different point source separations ($\theta$) ranging from 0.1 Rayleigh units to 1.2 Rayleigh units, resulting in 2400 trials each for our two-stage adaptive receiver and for direct imaging alone. For each one of these 4800 trials, the centroid ($\phi$) was randomly sampled from a uniform distribution across a domain of 100 SLM pixels (1.23 Rayleigh units). The two point sources for a given trial were therefore located at $\phi+\theta/2$ and $\phi-\theta/2$, which correspond to 1D pixel values on the SLM.
First, direct imaging data was acquired and used to estimate the centroid of the two point sources (Section 5A). We then use the centroid estimate as an input to our MPLC algorithm [20] to generate phase masks that sort PAD modes that are centered at the location of the centroid estimate in the object plane. This adaptive generation of a special-purpose set of phase masks for each trial effectively aligns our system to the scene and eliminates any need for physical manipulations. We then used the MPLC formed by the computed phase masks to perform the second-stage SPADE measurement. From this acquired data, we estimate the separation between the two point sources using maximum likelihood estimation as part of the data processing.
All the data in this experiment was taken using the photon counting mode in our EMCCD and the total number of photon arrivals for each trial was aimed to be 10,000. This means that each of the two stages detected 5,000 photons, on average, whereas the direct imaging measurement used for comparison detected all 10,000 photons. The laser power going into the system was calibrated accordingly while ensuring a low likelihood of multiple photons events. (See supplementary materials section 4). We conducted the experiment using a laser (coherent source), while for the experiment to correctly correlate the theory of super-resolution for two incoherent point sources, the light from the two point sources cannot interfere with each other. We therefore acquired data from only one point source at a time and then combined the two images computationally.